OpenStudy (anonymous):

You are going to invest $25, 000, part at 14% and part at 16%. What is the most that can be invested at 14% in order to make at least $3, 600 interest per year?

OpenStudy (ness9630):

If I'm not mistaken, you have to make two separte equations: \[\LARGE x+y=25000\] \[\LARGE 0.14x+0.16y=3600\]

OpenStudy (anonymous):

hmm.. i dont get it..

OpenStudy (ness9630):

Okay, so we are investing $25,000 right? Well we have to separate them evenly into two different accounts so we can earn an additional $3,600. The x+y just says that this much money plus this other half add up to the total $25,000.

OpenStudy (ness9630):

With those given equations we can find out how much x and y are.

OpenStudy (anonymous):

uh..

OpenStudy (ness9630):

Has your teacher taught you this yet? Or is completely new?

OpenStudy (anonymous):

completely new.. he just gave us activities

OpenStudy (ness9630):

Are you completely lost on what I'm saying? Or do you at least have an idea on what I'm saying.

OpenStudy (anonymous):

lost sir. im sorry.

OpenStudy (austinl):

You are going to invest $25, 000, part at 14% and part at 16%. What is the most that can be invested at 14% in order to make at least $3, 600 interest per year? Ok we have two different "values" essentially that we are going to invest. x at 14% y at 16% Your total investment will be the sum of those two values. And your total investment is 25,000 \(x+y=25,000\) Make sense? Now, the interest rates are different for x and y. We want the interest rates of each of them added to be 3,600. \(0.14x+0.16y=3,600\) Now we have a system of equations that can be solved. \(x+y=25,000\) \(0.14x+0.16y=3,600\) Have you covered how to solve these?

OpenStudy (anonymous):

0.14x+0.16y=3600?

OpenStudy (austinl):

I'm not sure what you are asking.

OpenStudy (ness9630):

That is just for the interest, remember the 14% and 16% interest?

OpenStudy (anonymous):

0.14x+0.16y=3600? im dont know what to do next..

OpenStudy (austinl):

No, with these problems, you have to look at both equations. \(x+y=25,000\\ 0.14x+0.16y=3,600\) What we can do is solve for y in the first equation, and plug it into the second equation. \(y=25,000-x\) \(0.14x+0.16(25,000-x)=3,600\) Then we can solve for x. Do you think you could solve it from there?

OpenStudy (ness9630):

Austin might take the spotlight now xD

OpenStudy (ness9630):

@MysteryOwl For now, just worry about solving the equation.

OpenStudy (anonymous):

let me try it :)

OpenStudy (anonymous):

is this correct? 0.14x + o.16 (25000-x)=3600 0.14x + 4000-0.16x=3600 -0.02x =3600-4000 -0.02x=-400 -0.02x/-0.02=-400/-0.02 x=200000

OpenStudy (ness9630):

Is that 20,000 or 200,000?

OpenStudy (anonymous):

200,000

OpenStudy (ness9630):

That's strange..

OpenStudy (anonymous):

why?

OpenStudy (ness9630):

How is it possible to put in $200,000 when you only have 20,000?

OpenStudy (anonymous):

:O something must be wrong in my solution.

OpenStudy (ness9630):

But anyways, forget that one, try solving for y, that one will make sense.

OpenStudy (ness9630):

I promise you, if we solve for y it will make so much more sense. :)

OpenStudy (anonymous):

okies xD let me try again :)

OpenStudy (anonymous):

y=5000

OpenStudy (ness9630):

Now that makes sense. So now that we have y, we can plug that back into the equation to solve for x. So now solve: x+5000=25000

OpenStudy (anonymous):

x=2000

OpenStudy (ness9630):

So 2,000+5,000=25,000?

OpenStudy (anonymous):

no xD its 20000

OpenStudy (ness9630):

There you go :) And that's pretty much it for solving.

OpenStudy (anonymous):

so the answer are x=20000 and y=5000.. are there no other solutions to make?

OpenStudy (ness9630):

You could of also checked your answers by graphing the systems:

OpenStudy (ness9630):

And nope, that's it.

OpenStudy (anonymous):

that was an awesome :) thanks..

OpenStudy (ness9630):

And sorry about the confusion earlier, as you go further into the chapter you will understand it more.